A veterinarian in the barangay wanted to determine the average number of pets. They got the following values.
1 5 2 2 4 1 2 2 4 3
a. Determine the 99% confidence interval for the population mean.
b. Determine the 95% confidence interval for the population mean
1
Expert's answer
2020-11-02T20:04:28-0500
"mean=\dfrac{1+5+2+2+4+1+2+2+4+3}{10}=2.6""s^2=\dfrac{\sum{(x_i-\bar{x})^2}}{n-1}=\dfrac{16.4}{9}"

"s=\sqrt{s^2}=\dfrac{2\sqrt{4.1}}{3}\approx1.3499"

a. The number of degrees of freedom are "df=10-1=9," and the significance level is "\alpha=0.01."

The critical t-value for "\alpha=0.01" and df = 9 degrees of freedom is "t_c=3.249823."

The 99% confidence for the population mean "\mu"  is computed using the following expression


"CI=(\bar{x}-t_c\times \dfrac{s}{\sqrt{n}}, \bar{x}+t_c\times \dfrac{s}{\sqrt{n}})="

"=(2.6-3.2498\times \dfrac{1.3499}{\sqrt{10}}, 2.6+3.2498\times \dfrac{1.3499}{\sqrt{10}})="

"=(1.2127, 3.9873)"

b. The number of degrees of freedom are "df=10-1=9," and the significance level is "\alpha=0.05."

The critical t-value for "\alpha=0.05" and df = 9 degrees of freedom is "t_c=2.262156."

The 95% confidence for the population mean "\mu"  is computed using the following expression


"CI=(\bar{x}-t_c\times \dfrac{s}{\sqrt{n}}, \bar{x}+t_c\times \dfrac{s}{\sqrt{n}})="

"=(2.6-2.262156\times \dfrac{1.3499}{\sqrt{10}}, 2.6+2.262156\times \dfrac{1.3499}{\sqrt{10}})="

"=(1.6343, 3.5657)"